(x)=|x²-x| differentiable at x=2. find its derivative

Grade:12th pass

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
f(x)=x^2-x for x>1
f'(x)=2x-1
f'(2)=4-1=3
So yes.
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty